OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -1 - 13t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (rational):

start by finding the derivative of s(t) ?

OpenStudy (anonymous):

i dont know how to find the derivative of -1-13 because there are no exponents or anything

OpenStudy (rational):

\[s(t) = -1-13t\] whats the derivative of \(-1\) with respect to \(t\) ?

OpenStudy (anonymous):

is it -8?

OpenStudy (rational):

nope derivative of a constant is 0

OpenStudy (anonymous):

ohh okay, so then what would the instantaneous velocity be?

OpenStudy (rational):

there is another term in s(t) whats the derivative of \(-13t\) ?

OpenStudy (rational):

if it helps notie that \(-13t\) is same as \(-13t^1\)

OpenStudy (anonymous):

-13t^0?

OpenStudy (rational):

yes! simplify

OpenStudy (anonymous):

-13?

OpenStudy (rational):

thats it! \[s(t) = -1-13t\] \[s'(t) = 0 - 13 = -13\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (rational):

yw!